OpenStudy (chihiroasleaf):

If \[A = \frac{1}{1+1}+\frac{1}{1+2}+\frac{1}{1+3}+...+\frac{1}{1+50}\] dan \[B = \frac{1}{1+1}+\frac{1}{1+\frac{1}{2}}+\frac{1}{1+\frac{1}{3}}+...+\frac{1}{1+\frac{1}{50}}\] then the value of \(5A^2+5B^2+10AB=...\)

mathslover (mathslover):

So, basically, we have to find : \(5(A+B)^2\) Right?

OpenStudy (chihiroasleaf):

yes.., but I have no idea how to simplify A and B

mathslover (mathslover):

Now, try this! A+ B

mathslover (mathslover):

Well, one minute. Writing LaTeX

OpenStudy (amistre64):

is the latex coding for yall?

mathslover (mathslover):

Yes!

OpenStudy (amistre64):

must be my computer then

OpenStudy (vishweshshrimali5):

Well, let @mathslover solve it... then I will show you a trick ;)

mathslover (mathslover):

I will represent terms of A in red and that of B in blue \(\color{red}{A} + \color{blue}{B} = \left( \color{red}{\cfrac{1}{1+1}} + \color{blue}{\cfrac{1}{1+1}} \right) + \left( \color{Red}{\cfrac{1}{1+2}} + \color{blue}{\cfrac{1}{1+1/2}} \right) + ... + \left( \color{red}{\cfrac{1}{1+50}} + \color{blue}{\cfrac{1}{1+1/50}} \right) \)

OpenStudy (vishweshshrimali5):

Well, looks like that he caught it :D

OpenStudy (chihiroasleaf):

ahhh..., I get it... :))

OpenStudy (chihiroasleaf):

A+B = 51 ?

mathslover (mathslover):

\(\color{red}{A} + \color{blue}{B} = \left( \color{red}{\cfrac{1}{1+1}} + \color{blue}{\cfrac{1}{1+1}} \right) + \left( \color{Red}{\cfrac{1}{1+2}} + \color{blue}{\cfrac{1}{1+1/2}} \right) + ... \\ + \left( \color{red}{\cfrac{1}{1+50}} + \color{blue}{\cfrac{1}{1+1/50}} \right) \)

mathslover (mathslover):

Now, we have : \(A + B = 1 + 1 + ... + 1\) Clearly it is 51 :)

mathslover (mathslover):

This is how I came up with that : \(\cfrac{1}{1+1} + \cfrac{1}{1+1} = 1\) \(\cfrac{1}{1+2} + \cfrac{1}{1 +1/2} = \cfrac{1}{3} + \cfrac{2}{3} = 1\) And so on.. So you get 1 + 1 + .. + 1 (50 times) = 50 (Sorry for the mistake earlier, it is 50 not 51)

mathslover (mathslover):

@vishweshshrimali5 - What was the trick you were talking about? Am curious to know about it

OpenStudy (chihiroasleaf):

ahhh..., I got the same mistake.., yes.., it's 50 so, \(5(50)^2 \)

OpenStudy (vishweshshrimali5):

Well... I really like using \(\Sigma\) So: \[\large{A = \Sigma_{i = 1}^{50} \cfrac{1}{1+i}}\] \[\large{B = \Sigma_{i = 1}^{50} \cfrac{1}{1+\cfrac{1}{i}}} = \Sigma \cfrac{i}{1+i}\] \[\large{A+B = \Sigma_{i=1}^{50} 1 = 50}\] Thus: \(\large{5(A+B)^2 = 5 *(50)^2}\)